# Bilayer graphene in periodic and quasiperiodic magnetic superlattices

### v1

@@ Abstract
Starting from the effective Hamiltonian arising from the tight binding model, we study the behaviour of low-lying excitations for bilayer graphene placed in periodic external magnetic fields by using irreducible second order supersymmetry transformations. The coupled system of equations describing these excitations is reduced to a pair of periodic Schrödinger Hamiltonians intertwined by a second order differential operator. The direct implementation of more general second-order supersymmetry transformations allows to create nonsingular Schrödinger potentials with periodicity defects and bound states embedded in the forbidden bands, which turn out to be associated to quasiperiodic magnetic superlattices.
@@ 1 Introduction
Due to its Dirac-like low lying excitations, graphene displays a number of unusual transport properties, which results in the pseudospin 1/2 of the low-lying modes, their linear dispersion relation, and the vanishing density of states at the Dirac points. In contrast, the interlayer interaction in bilayer graphene with regular AB stacking changes the linear dispersion of monolayer graphene into a quadratic one, where an electron behaves as a massive particle [15]. In order to describe the situation when magnetic (or pseudomagnetic) fields are applied to monolayer graphene, the first-order supersymmetric quantum mechanics (SUSY QM) is the simplest solution approach [16, 17, 18, 19, 21, 20, 22, 23, 24] (see also [25, 26, 27, 28, 29, 30, 31]). On the other hand, the second-order SUSY QM has proved useful to study the charge carriers behaviour of bilayer graphene in external magnetic fields [32, 33]. In addition, the equivalence between Maxwell’s and Dirac equation has been used to understand the electromagnetic spin and orbital angular momentum, and examine the relationship between interface states and topology [34, 35, 36, 37].
Recently, the first order SUSY QM has been used to study the electronic behaviour in superlattices, i.e, when monolayer graphene is placed in periodic magnetic fields [38]. The arising of two opposite Darboux displacement breaks the periodicity in the superlattice, but it can be shown that asymptotically the periodicity is recovered. Moreover, the use of such displacements turns out to be fundamental for understanding the phenomenon related to defects in superlattices.
@@
In the present work we aim to study the behaviour of low-lying excitations of bilayer graphene placed in a periodic magnetic field, and its physical implications, by using second order supersymmetric quantum mechanics. We will explore as well more general second-order supersymmetry transformations producing periodicity defects, which will be associated naturally to quasiperiodic magnetic superlatticces for bilayer graphene. In order to do that, the second section is devoted to some general considerations about the tight-binding model for bilayer graphene, its physical assumptions and the two coupled equations we will arrive for our system. Then, in section 3 a brief overview of the second order SUSY QM will be given, as well as the construction of the intertwining operators when a periodic potential is present. The way to implement the method for generating quasiperiodic partner potentials and the corresponding superpotential will be also discussed. In section 4 we will generate the quasiperiodic magnetic superlattices associated to a physically meaningful quantum problem. Finally, in section 5 we will present our conclusions and the perspectives of this work.
@@ 2 Tight-binding model for bilayer graphene
Since we want to study the case when a periodic magnetic field is applied, $\eta(x)$ and $\gamma(x)$ in equations (4) and (5) are supposed to be complex valued piecewise continuous periodic functions, both with the same period. Note that an iterative first order SUSY treatment could be used, such that the second order differential intertwining operators (4) and (5) will be factorized as a product of two in general different first order intertwining operators. These assumptions require to modify the effective Hamiltonian by an additional term, that could be physically interpreted as the result of a trigonal warping effect or a varying external potential. This is expressed mathematically as follows
@@
We have chosen the index notation slightly different from [32, 33], for reasons that will be evident below. As can be seen, in order to study the effect of periodic magnetic fields in bilayer graphene, the second order supersymmetric quantum mechanics can be used, as when studying the charge carriers behaviour of bilayer graphene in non-periodic external magnetic fields. Before doing that, however, a brief overview of the second order supersymmetric quantum mechanics will be given in the following section.
@@ 3 Second order supersymmetric quantum mechanics
Let us assume that $v_{1}(x)$ and $v_{2}(x)$ are two solutions of the stationary Schrödinger equation for the Hamiltonian $H_{0}$ associated to the factorization energies $\epsilon_{1}$ and $\epsilon_{2}$, respectively, i.e., $H_{0}v_{i}=\epsilon_{i}v_{i}$, $i=1,2$, where
@@
We will assume as well that both Hamiltonians (12) and (14) have discrete spectra. The so-called SUSY partner potentials $V_{0}(x)$ and $V_{2}(x)$ can be expressed in terms of the functions $\eta(x)$ and $\gamma(x)$ characterizing the interwining operator $L_{2}^{-}$. After some work, the following expressions are obtained:
@@
${V}_{2}(x)=V_{0}(x)+2\eta^{{}^{\prime}}(x)=V_{0}(x)-2\dfrac{d^{2}}{dx^{2}}\left(\ln W_{12}(x)\right).$ (19)
It will be free of singularities whenever $W_{12}(x)$ is nodeless. The eigenfunctions of $H_{0}$ and ${H}_{2}$, $\{\psi_{i}^{(0)}(x),{\psi}_{i}^{(2)}(x),i=0,1,\dots\}$, are connected by the operators $L_{2}^{-}$, $L_{2}^{+}$ as follows:
@@
Depending on whether they can be normalized or not, such factorization energies can be either included or not in the spectrum of $H_{2}$. Therefore, the Hamiltonians $H_{0}$ and $H_{2}$ are isospectral, with the possible exception of $\epsilon_{1}$ and $\epsilon_{2}$, which could be in Sp$(H_{2})$. From now on we will assume that $V_{0}(x)$ is a given initial periodic potential of period $T$, $V_{0}(x+T)=V_{0}(x)$. In this case there is no need to consider the range of the variable $x$ as the whole real line, but just a given period. According to the Bloch-Floquet theory, the physical wavefunctions $\psi(x)$ are quasi-periodic:
@@ 3.1 Lamé equation: some remarks
The stationary Schrödinger equation for the Hamiltonian (12) with the periodic potential given by expression (23) is the Jacobi version of the Lamé equation, where $\text{sn}(x|m)$ is the Jacobi elliptic function in Glaisher notation with modular parameter $m$. When $m\in(0,1)$ the function $\text{sn}^{2}(x|m)$ has real period $2K=2K(m)$ and an imaginary period $2iK^{\prime}=2iK(1-m)$, with $K(m)$ being the first complete elliptic integral. In addition, if $x$ is restricted to the real axis and $m$ and $n$ are real, the Lamé equation becomes a real domain Schrödinger equation with a periodic potential, i.e., a Hill’s equation.
@@
with $\xi=\pm 1$ being a Floquet multiplier. Due to the fact that the $H_{0}$ in Eq. (12) with the $V_{0}(x)$ given by expression (23) is a Lamé operator of Hill type, we know from the oscillation theorem that there exists a sequence of real numbers (in general infinite)
@@
where the band edge energies $E_{j}$ and $E_{j^{\prime}}$ correspond to $2K$-periodic solutions for $j$ even and $2K$-antiperiodic solutions for $j$ odd. The physical energies lie in allowed bands, which are intervals delimited by energies corresponding to the two values of $\xi$, i.e., to periodic and anti-periodic Bloch solutions. These allowed bands form a sequence, where $E_{0}$ is the first periodic eigenvalue if we come from $\infty$- ∞, followed by alternating pairs of anti-periodic and periodic eigenvalues (each pair may be coincident). The allowed energy bands are $[E_{0},E_{1}]$, $[E_{1^{\prime}},E_{2}],...,[E_{j^{\prime}},E_{j+1}],...$; consistently, the forbidden energy bands (energy gaps) are $(E_{1},E_{1^{\prime}})$, $(E_{2},E_{2^{\prime}}),...,(E_{j},E_{j^{\prime}}),...$ The band edge energies are usually called the discrete spectrum of the periodic potential. The corresponding eigenfunctions $\{\psi_{0}^{(0)},\psi_{1}^{(0)},\psi_{1^{\prime}}^{(0)},\psi_{2}^{(0)},\psi_{2^{\prime}}^{(0)},\dots\}$ have periods $T$, $2T$, $2T$, $T$, $T$, $\dots$, and their respective number of nodes in a period $T$ is $0,1,1,2,2,\dots$.
@@ 4 Bilayer graphene in periodic magnetic fields
This happens since it is known that the self-isospectrality condition is fulfilled if $\psi_{j}^{(0)}(x)\psi_{j}^{(0)}(x+T/2)\propto\psi_{j^{\prime}}^{(0)}(x)\psi_{j^{\prime}}^{(0)}(x+T/2)$ [60, 61]. In our case, choosing the band edge eigenfunctions $\psi_{1}^{(0)}$ and $\psi_{1^{\prime}}^{(0)}$ given in equations (39) and (40) we will have
@@
which is the condition for $V_{2}(x)$ to be physically equivalent (self-isospectral) to $V_{0}(x)$. Thus, there is no new potential obtained for $n=1$ if band edge eigenfunctions are used. However, for the Lamé potential (23) with $n=2,3,...$, the self-isospectrality condition is not longer fulfilled, and thus new potentials could be obtained [52, 53, 54].
Let us point out that the general Darboux-Crum transformations have been applied to the single gap periodic Lamé potential to generate an arbitrary countable set of bound states in its two forbidden bands, the infinite lowest and the finite intermediate one. As a consequence, nonperiodic single gap potentials arise, which contain two essentially different types of soliton deffects in the periodic background [56, 57]. As it was already shown, these defects appear in a natural way in carbon-based crystals such as monolayer graphene [38] and, as we will see below, for bilayer graphene. In order to do that, we need to use as seed solutions linear combinations of the two generalized Bloch functions belonging to the same spectral gap in order to obtain the new potential. In this way, the second-order transformation will produce a quasi-periodic magnetic superlattice supporting two bound states.
Let us start from the Lamé magnetic superlattice; the seed solutions we will consider are $v_{j}=u_{1}(x,\epsilon_{j})+\eta_{j}u_{2}(x,\epsilon_{j})$, $j=1,2$ with $\epsilon_{2}<\epsilon_{1}<0$, which fulfill the Lamé equation. The two Bloch functions $u_{j}$ take the form
@@
where $j=1,2$, $x_{0}$ is a fixed point in $[0,T=2K]$, $\omega=K$ and $\omega=iK^{\prime}$ are the real and imaginary half-periods of the Jacobi elliptic functions, $\sigma$ and $\zeta$ are the nonelliptic Weierstrass functions, while the factorization energy $\epsilon_{j}$ and the displacement $\delta_{j}$ are related by
@@
A short hand notation for $u(x,\epsilon_{j})$ in the Wronskians will be employed
@@
In particular, for the limit case $\eta_{1}\to 0$ and $\eta_{2}\to 0$ the Wronskian (46) becomes (47), which is consistent with [61]. In the general case, the explicit expression for the magnetic field with impurities is obtained by replacing Eq. (44) in (7):
@@
with asymptotic behaviour given by $(x\to\pm\infty)$
@@
As it was expected, the expressions for the magnetic field with two added impurities are similar to the ones obtained for the case of a single displacement with an added impurity [38]. It is straightforward to find the expressions for the added bound states in the forbidden band $(E_{1},E_{1^{\prime}})$ for bilayer graphene, which arise from equation (21) using linear combinations $v_{1}$ and $v_{2}$ of the Bloch functions (see a plot in Fig. 1(a)). Let us stress that, the implementation of the second order supersymmetry transformation for bilayer graphene is mandatory, unlike what happens for monolayer graphene in which an additional SUSY transformation could be applied to go deeper in the study of impurity effects, but it is not strictly required.
Figure 1: (a) Bound states created at the energy $\alpha_{1}=0.9$ and at $\alpha_{2}=0.8$ (b) for $m=0.5$.
@@
Throughout this work we have restricted ourselves to study the effect of the Lamé periodic potentials in bilayer graphene; nevertheless, the basic ideas of this work could be applied to more general potentials, e.g., to the associated Lamé potentials [54, 58, 59]. We are confident that the basic idea of introducing defects in the periodic magnetic superlattice for bilayer graphene by adding bound states in the forbidden bands, can be well understood through one of the simplest periodic potentials at hand.

### v2

@@ Abstract
Starting from the effective Hamiltonian arising from the tight binding model, we study the behaviour of low-lying excitations for bilayer graphene placed in periodic external magnetic fields by using irreducible second order supersymmetry transformations. The coupled system of equations describing these excitations is reduced to a pair of periodic Schrödinger Hamiltonians intertwined by a second order differential operator. The direct implementation of more general second-order supersymmetry transformations allows to create nonsingular Schrödinger potentials with periodicity defects and bound states embedded in the forbidden bands, which turn out to be associated to quasiperiodic magnetic superlattices. Applications in quantum metamaterials stem from the ability to engineer and control such bound states which could lead to a fast development of the subject in the near future.
@@ 1 Introduction
Due to its Dirac-like low lying excitations, graphene displays a number of unusual transport properties, which results in the pseudospin 1/2 of the low-lying modes, their linear dispersion relation, and the vanishing density of states at the Dirac points. In contrast, the interlayer interaction in bilayer graphene with regular AB stacking changes the linear dispersion of monolayer graphene into a quadratic one, where an electron behaves as a massive particle [15]. In order to describe the situation when magnetic (or pseudomagnetic) fields are applied to monolayer graphene, the first-order supersymmetric quantum mechanics (SUSY QM) is the simplest solution approach [16, 17, 18, 19, 21, 20, 22, 23, 24] (see also [25, 26, 27, 28, 29, 30, 31]). On the other hand, the second-order SUSY QM has proved useful to study the charge carriers behaviour of bilayer graphene in external magnetic fields [32, 33]. In addition, the equivalence between Maxwell’s and Dirac equation has been used to understand the electromagnetic spin and orbital angular momentum, and examine the relationship between interface states and topology [34, 35, 36, 37, 38, 39].
Recently, there has been a growing interest in studying and characterizing transport properties in superlattices composed of stacked van der Waals heterostructures, largely motivated by magic-angle twisted graphene [40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. In this direction, numerous experimental investigations have been realized on dual-gated hexagonal boron nitride and bilayer graphene superlattices, revealing first-order transitions between two insulating states under a perpendicular magnetic field [50, 51, 52]. On the theoretical side, first-order SUSY QM has been used to study the electronic behavior in superlattices featuring monolayer graphene subjected to periodic magnetic fields [53]. It has been observed that the arising of two opposite Darboux displacements disrupts the superlattice’s periodicity, although asymptotically the periodicity is recovered. Moreover, the use of such displacements turns out to be fundamental for understanding phenomena related to defects in superlattices.
@@
In the present work we aim to study the behaviour of low-lying excitations of bilayer graphene placed in a periodic magnetic field, and its physical implications, by using second order supersymmetric quantum mechanics. We will explore as well more general second-order supersymmetry transformations producing periodicity defects, which will be associated naturally to quasiperiodic magnetic superlatticces for bilayer graphene. In order to do that, the second section of this paper is devoted to some general considerations about the tight-binding model for bilayer graphene, its physical assumptions and the two coupled equations we will arrive for our system. Then, in section 3 a brief overview of the second order SUSY QM will be given, as well as the construction of the intertwining operators when a periodic potential is present. The way to implement the method for generating quasiperiodic partner potentials and the corresponding superpotential will be also discussed. In section 4 we will generate the quasiperiodic magnetic superlattices associated to a physically meaningful quantum problem. Finally, in section 5 we will present our conclusions and the perspectives of this work.
@@ 2 Tight-binding model for bilayer graphene
In general, after generating the vector potential $A(x)$ through a second-order SUSY transformation, the amplitude of the magnetic field $B(x)$ (or pseudo-magnetic field) given by expression (7) can be directly obtained. Depending on the specific physical context under investigation, different forms of $A(x)$ can be proposed, each associated with the superpotential as it will be shown in the following section. Since we want to study the case when a periodic magnetic field is applied, $\eta(x)$ and $\gamma(x)$ in equations (4) and (5) are supposed to be complex valued piecewise continuous periodic functions, both with the same period. Note that an iterative first order SUSY treatment could be used, such that the second order differential intertwining operators (4) and (5) will be factorized as a product of two in general different first order intertwining operators. These assumptions require to modify the effective Hamiltonian by an additional term, that could be physically interpreted as the result of a trigonal warping effect or a varying external potential. This is expressed mathematically as follows
@@
We have chosen the index notation slightly different from [32, 33], for reasons that will be evident below. As can be seen, in order to study the effect of periodic magnetic fields on bilayer graphene, the second order supersymmetric quantum mechanics can be used, as when studying the charge carriers behaviour of bilayer graphene in non-periodic external magnetic fields. Before doing that, however, a brief overview of the second order supersymmetric quantum mechanics will be given in the following section.
@@ 3 Second order supersymmetric quantum mechanics
Now, we present the algorithm of second order supersymmetric quantum mechanics for showing its explicit relation with bilayer graphene placed in periodic magnetic fields [54]. First, let us assume that $v_{1}(x)$ and $v_{2}(x)$ are two solutions of the stationary Schrödinger equation for the Hamiltonian $H_{0}$ associated to the factorization energies $\epsilon_{1}$ and $\epsilon_{2}$, respectively, i.e., $H_{0}v_{i}=\epsilon_{i}v_{i}$, $i=1,2$, where
@@
We will assume as well that both Hamiltonians (12) and (14) have discrete spectra. The so-called SUSY partner potentials $V_{0}(x)$ and $V_{2}(x)$ can be expressed in terms of the functions $\eta(x)$ and $\gamma(x)$ characterizing the interwining operator $L_{2}^{-}$. After some work, in the first place it is obtained that
@@
${V}_{2}(x)=V_{0}(x)+2\eta^{{}^{\prime}}(x)=V_{0}(x)-2\dfrac{d^{2}}{dx^{2}}\left(\ln W_{12}(x)\right),$ (19)
which will be free of singularities whenever $W_{12}(x)$ is nodeless. The eigenfunctions of $H_{0}$ and ${H}_{2}$, $\{\psi_{i}^{(0)}(x),$ ${\psi}_{i}^{(2)}(x),i=0,1,\dots\}$, are connected by the operators $L_{2}^{-}$, $L_{2}^{+}$ as follows:
@@
Depending on whether they can be normalized or not, such factorization energies must be either included or not in the spectrum of $H_{2}$. Therefore, the Hamiltonians $H_{0}$ and $H_{2}$ are isospectral, with the possible exception of $\epsilon_{1}$ and $\epsilon_{2}$, which could be in Sp$(H_{2})$. From now on we will assume that $V_{0}(x)$ is a given initial periodic potential of period $T$, $V_{0}(x+T)=V_{0}(x)$. In this case there is no need to consider the range of the variable $x$ as the whole real line, but just a given period. According to the Bloch-Floquet theory, the physical wavefunctions $\psi(x)$ are quasi-periodic:
@@ 3.1 Lamé equation: some remarks
The stationary Schrödinger equation for the Hamiltonian (12) with the periodic potential (23) is the Jacobi version of the Lamé equation, where $\text{sn}(x|m)$ is the Jacobi elliptic function in Glaisher notation with modular parameter $m$. When $m\in(0,1)$ the function $\text{sn}^{2}(x|m)$ has real period $2K=2K(m)$ and an imaginary period $2iK^{\prime}=2iK(1-m)$, with $K(m)$ being the first complete elliptic integral. In addition, if $x$ is restricted to the real axis and $m$ and $n$ are real, the Lamé equation becomes a real domain Schrödinger equation with a periodic potential, i.e., a Hill’s equation.
@@
with $\xi=\pm 1$ being a Floquet multiplier. Due to the fact that $H_{0}$ in Eq. (12) with $V_{0}(x)$ given by expression (23) is a Lamé operator of Hill type, we know from the oscillation theorem that there exists a sequence of real numbers (in general infinite)
@@
where the band edge energies $E_{j}$ and $E_{j^{\prime}}$ correspond to $2K$-periodic solutions for $j$ even and $2K$-antiperiodic solutions for $j$ odd. The physical energies lie in allowed bands, which are intervals delimited by energies corresponding to the two values of $\xi$, i.e., to periodic and anti-periodic Bloch solutions. These allowed bands form a sequence, where $E_{0}$ is the first periodic eigenvalue if we come from $\infty$- ∞, followed by alternating pairs of anti-periodic and periodic eigenvalues (each pair might be coincident). The allowed energy bands are $[E_{0},E_{1}]$, $[E_{1^{\prime}},E_{2}],...,[E_{j^{\prime}},E_{j+1}],...$; consistently, the forbidden energy bands (energy gaps) are $(E_{1},E_{1^{\prime}})$, $(E_{2},E_{2^{\prime}}),...,(E_{j},E_{j^{\prime}}),...$ The band edge energies are usually called the discrete spectrum of the periodic potential. The corresponding eigenfunctions $\{\psi_{0}^{(0)},\psi_{1}^{(0)},\psi_{1^{\prime}}^{(0)},\psi_{2}^{(0)},\psi_{2^{\prime}}^{(0)},\dots\}$ have periods $T$, $2T$, $2T$, $T$, $T$, $\dots$, and their respective number of nodes in a period $T$ is $0,1,1,2,2,\dots$.
@@ 4 Bilayer graphene in periodic magnetic fields
This happens since it is known that a self-isospectral condition is fulfilled if $\psi_{j}^{(0)}(x)\psi_{j}^{(0)}(x+T/2)\propto\psi_{j^{\prime}}^{(0)}(x)\psi_{j^{\prime}}^{(0)}(x+T/2)$ [68, 69]. In our case, choosing the band edge eigenfunctions $\psi_{1}^{(0)}$ and $\psi_{1^{\prime}}^{(0)}$ given in equations (39) and (40) we will have
@@
which is the condition for $V_{2}(x)$ to be physically equivalent (self-isospectral) to $V_{0}(x)$. Thus, there is no new potential obtained for $n=1$ if band edge eigenfunctions are used. However, for the Lamé potential (23) with $n=2,3,...$, the self-isospectral condition is not longer fulfilled, and thus new potentials could be obtained [70, 71, 72, 73].
Let us point out that the general Darboux-Crum transformations have been applied to the single gap periodic Lamé potential to generate an arbitrary countable set of bound states in its two forbidden bands, the infinite lowest and the finite intermediate one. As a consequence, nonperiodic single gap potentials arise, which contain two essentially different types of soliton defects in the periodic background [74, 75, 76]. As it was already shown, these defects appear in a natural way in carbon-based crystals such as monolayer graphene [53] and, as we will see below, for bilayer graphene. In order to do that, we need to use as seed solutions linear combinations of the two generalized Bloch functions belonging to the same spectral gap in order to obtain the new potential. In this way, the second-order transformation will produce a quasi-periodic magnetic superlattice supporting two bound states.
Let us start from the Lamé magnetic superlattice; the seed solutions we will consider, which fulfill the Lamé equation, are $v_{j}=u_{1}(x,\epsilon_{j})+\eta_{j}u_{2}(x,\epsilon_{j})$, $j=1,2$ with $\epsilon_{2}<\epsilon_{1}<0$. The two Bloch functions $u_{j}$ take the form
@@
where $j=1,2$, $x_{0}$ is a fixed point in $[0,T=2K]$, $\omega=K$ and $\omega^{\prime}=iK^{\prime}$ are the real and imaginary half-periods of the Jacobi elliptic functions, $\sigma$ and $\zeta$ are the nonelliptic Weierstrass functions, while the factorization energy $\epsilon_{j}$ and the complex displacement $\delta_{j}$ are related by
@@
A short hand notation for $u(x,\epsilon_{j})$ in the Wronskians will be employed from now on,
@@
with asymptotic behaviour for $x\to\pm\infty$ given by
@@
Expression (48) represents a Lamé magnetic superlattice with two additional impurities, analogous to the case obtained for monolayer graphene with a single displacement and an added impurity [53]. It is straightforward to find the expressions for the added bound states in the forbidden band $(E_{1},E_{1^{\prime}})$ for bilayer graphene, which arise from equation (21) using the linear combinations $v_{1}$ and $v_{2}$ of the Bloch functions (see a plot in Fig. 1). Also, we can generate the partner potentials $V_{0}(x)$ and $V_{2}(x)$ given by expressions (17) and (18) for our linear combination of Bloch functions, once the Wronskian (47) is found for the arbitrary displacements $\delta_{j}$. Let us stress that the implementation of the second order supersymmetry transformation for bilayer graphene is mandatory, unlike what happens for monolayer graphene where an additional first order SUSY transformation could be applied to go deeper in the study of impurity effects, but it is not strictly required.
Figure 1: Bound states created at the energy $\alpha_{1}=0.9$ (a) and at $\alpha_{2}=0.8$ (b) for $m=0.5$.
@@
Throughout this work we have restricted ourselves to study the effect of the Lamé periodic potentials in bilayer graphene; nevertheless, the basic ideas of this work could be applied to more general potentials, e.g., to the associated Lamé potentials [77, 78]. We are confident that the basic idea of introducing defects in the periodic magnetic superlattice for bilayer graphene, by adding bound states in the forbidden bands, can be well understood through the simplest periodic potentials at hand.